OpenStudy (anonymous):

Find the distance, in meters, a particle travels in its first 10 seconds of travel, if it moves according to the velocity equation v(t)= 49 − 9.8t (in meters/sec). A: −49 B: 0 C: 245 D: 441

OpenStudy (anonymous):

@primeralph can u plz help? :)

OpenStudy (anonymous):

@campbell_st

OpenStudy (xapproachesinfinity):

i don't really see any question

OpenStudy (xapproachesinfinity):

you just stated the info

OpenStudy (xapproachesinfinity):

o o o i take it back lol you did have the question there

OpenStudy (anonymous):

its ok :)

OpenStudy (xapproachesinfinity):

okay let's see what do we have here

OpenStudy (xapproachesinfinity):

ok so \[\Delta x=49t-4.9t^2 \] evaluate at t=10

OpenStudy (anonymous):

0

OpenStudy (anonymous):

triangle x = 0

OpenStudy (xapproachesinfinity):

this is one of the position equations you should have something similar to this in your notes

OpenStudy (xapproachesinfinity):

what are you talking about?

OpenStudy (anonymous):

i plugged in t= 10 into the deltax = 49t-4.9t^2

OpenStudy (xapproachesinfinity):

this is what i meant \[\Delta x =4.9(10)-4.9(10)^2\]

OpenStudy (xapproachesinfinity):

it's not zero

OpenStudy (anonymous):

49-490=-441

OpenStudy (anonymous):

So it would be: D, 441! thank u!

OpenStudy (xapproachesinfinity):

hmm the answer you are given is positive the one we found is negative something is off

OpenStudy (xapproachesinfinity):

was you equation \[v(t)=4.9-9.8t\]

OpenStudy (xapproachesinfinity):

your?

OpenStudy (anonymous):

v(t)= 49 − 9.8t (in meters/sec).

OpenStudy (xapproachesinfinity):

i can't see what you wrote lol

OpenStudy (anonymous):

its v(t)=49-9.8t

OpenStudy (anonymous):

no decimal between 4 & 9 :)

OpenStudy (xapproachesinfinity):

you sure?

OpenStudy (anonymous):

100%

OpenStudy (xapproachesinfinity):

okay then you have \[\Delta x=0\]

OpenStudy (xapproachesinfinity):

i see why you said that from the start haha

OpenStudy (anonymous):

perfect :) thank u

OpenStudy (anonymous):

i have 1 more question do u mind helping me with it?

OpenStudy (xapproachesinfinity):

np!

OpenStudy (xapproachesinfinity):

let's see

OpenStudy (anonymous):

For an object whose velocity in ft/sec is given by v(t) = −3t^2 + 5, what is its displacement, in feet, on the interval t = 0 to t = 2 secs? A: 6.607 B: 2 C: −2.303 D: 2.303

OpenStudy (anonymous):

I know its not A, 6.607 I think its C, -2.303

OpenStudy (xapproachesinfinity):

what did you do

OpenStudy (xapproachesinfinity):

did you compute the integral

OpenStudy (xapproachesinfinity):

that should gave the desplacement

OpenStudy (primeralph):

Just integrate.

OpenStudy (anonymous):

can u please walk me through the steps

OpenStudy (xapproachesinfinity):

do you know about integrals you are doing physics I yes?

OpenStudy (anonymous):

calc 1 :)

OpenStudy (xapproachesinfinity):

well you should know how to integrate unless you just started calc1

OpenStudy (xapproachesinfinity):

at any rate let's do it \[\int_{0}^{2}(-3t^2+5 )dx=\left[-t^3+5t\right]_{0}^{2}=-(2)^3+5(2)+0-0\]

OpenStudy (anonymous):

-8+10+0-0=2 :)

OpenStudy (anonymous):

brilliant thank u

OpenStudy (xapproachesinfinity):

that's the use of th fundamental theorem of calc

OpenStudy (xapproachesinfinity):

:) np!

OpenStudy (anonymous):

i got this 1 wrong :( Find the distance, in meters, a particle travels in its first 10 seconds of travel, if it moves according to the velocity equation v(t)= 49 − 9.8t (in meters/sec).

OpenStudy (xapproachesinfinity):

i don't think that's 49 should be 4.9 i guess

OpenStudy (xapproachesinfinity):

you can use the integral once more with that as well

OpenStudy (anonymous):

got it :) thank u

OpenStudy (anonymous):

For an object whose velocity in ft/sec is given by v(t) = −t2 + 6, what is its displacement, in feet, on the interval t = 0 to t = 3 secs? 9.000 −9.000 10.596 −3.00

OpenStudy (xapproachesinfinity):

they all the same problems my friend just follow the same procedure to answer them

OpenStudy (xapproachesinfinity):

just integrate again from 0 to 3

OpenStudy (anonymous):

can you please walk me through those steps again

OpenStudy (anonymous):

i know its either A,9.000 or C,10.596

OpenStudy (mathmate):

@mathrulezz The first question about the particle requires you to find the distance travelled, not displacement. So it is not displacement=\(x_{final} - x_{initial}\) It is total distance travelled. See diagrams below:  So the particle has travelled to x=122.5 and back, for a total \(distance\) of 245.